OpenStudy (anonymous):

Solve for x. 6 – 3(2x – 4) = 2(1 + x) A.–2 B.0 C.–1 D.2

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

6-3(2x-4)=2(1+x) 6-6x+12=2+2x -6x+18=2+2x -8x=-16 x=2 choice D